# Learning to Optimize for Mixed-Integer Non-linear Programming

### v4

@@ Appendix D Neural Network Structure and Hyperparameters
The constraint penalty weight $\lambda$ was set to $200$ for convex quadratic problems, $100$ for simple non-convex problems and Rosenbrock problems. All networks were trained using the AdamW optimizer with a learning rate of $10^{-3}$ and a batch size of $64$ over $200$ epochs. Early stopping was applied based on validation performance to ensure convergence without overfitting.
@@ Appendix E MINLP Problem Setup and Parameter Sampling
For the convex quadratic problems, $\bm{Q}$ is a diagonal matrix with entries drawn from a uniform distribution over $[0,0.01]$ to guarantee the convexity of the problem. The $\bm{p}$ was generated with entries drawn from a uniform distribution over $[0,0.1]$. The matrix $\bm{A}$ was drawn from a normal distribution with a standard deviation of $0.1$. The parameter $\bm{b}\in\mathbb{R}^{m}$, the right-hand side vector of the inequality constraints, is sampled from a uniform distribution over the range $[-1,1]$.
@@
For the simple non-convex problems, the setup remains consistent with that of the convex quadratic problems. However, an additional parameter $\bm{d}\in\mathbb{R}^{m}$ is introduced, with each element independently sampled from a uniform distribution over $[-0.5,0.5]$. The parameter $\bm{d}$ modifies the constraint matrix $\bm{A}$ by adding $\bm{d}$ to its first column and subtracting $\bm{d}$ from its second column.
@@
For the Ronsenbrock Problems, the vectors $\bm{p}\in\mathbb{R}^{n}$ and $\bm{q}\in\mathbb{R}^{n}$ are fixed for each instance and generated from a standard normal distribution. The parameter $b$ is uniformly distributed over $[1,8]$ for each instance, and the parameter $\mathbf{a}\in\mathbb{R}^{n}$ represents a vector where elements drawn independently from a uniform distribution over $[0.5,4.5]$. The parameters $b$ and $\mathbf{a}$ influence the shape of the feasible region and the landscape of the objective function, serving as input features to the neural network.
@@ Appendix F Ablation Study
The results of the ablation experiments, summarized in Table 2, demonstrate the importance of the correction layers $\varphi_{\Theta_{2}}$ in improving both solution quality and feasibility. RL shows a significant drop in feasibility rates, highlighting the importance of incorporating learnable corrective adjustments during training. Similarly, while RS benefits from differentiability via STE, the lack of learnable correction limits its performance compared to RC and LT.
@@
RL Obj Mean $0.567$ $-2.325$ $-4.576$ $-13.932$ $-16.734$ $-35.562$ $-88.217$ $-170.738$
Obj Median $0.405$ $-2.654$ $-4.606$ $-13.952$ $-16.803$ $-35.545$ $-88.217$ $-170.738$
% Infeasible $24\%$ $74\%$ $23\%$ $33\%$ $62\%$ $73\%$ $100\%$ $100\%$
Time (Sec) $0.0004$ $0.0005$ $0.0004$ $0.0004$ $0.0006$ $0.0005$ $0.0005$ $0.0011$
RS Obj Mean $1.024$ $-0.973$ $-3.677$ $-11.542$ $-10.092$ $-22.366$ $-53.036$ $-105.194$
Obj Median $0.880$ $-1.269$ $-3.706$ $-11.525$ $-10.073$ $-22.483$ $-53.036$ $-105.194$
@@
RL Obj Mean $0.241$ $0.595$ $-0.138$ $-0.629$ $-1.581$ $-4.196$ $-11.531$ $-23.639$
Obj Median $0.090$ $0.484$ $-0.148$ $-0.655$ $-1.554$ $-4.196$ $-11.531$ $-23.639$
% Infeasible $28\%$ $46\%$ $13\%$ $49\%$ $85\%$ $100\%$ $100\%$ $100\%$
Time (Sec) $0.0005$ $0.0005$ $0.0005$ $0.0005$ $0.0006$ $0.0006$ $0.0006$ $0.0013$
RS Obj Mean $0.453$ $1.398$ $0.292$ $1.734$ $2.849$ $4.921$ $9.511$ $25.358$
Obj Median $0.328$ $0.983$ $0.284$ $1.736$ $2.841$ $4.907$ $9.511$ $25.358$
@@
RL Obj Mean $58.34$ $63.70$ $605.88$ $6221.75$ $68364.18$
Obj Median $58.00$ $61.95$ $609.00$ $5949.55$ $69087.00$
% Infeasible $86\%$ $36\%$ $44\%$ $28\%$ $31\%$
Time (Sec) $0.0006$ $0.0005$ $0.0005$ $0.0008$ $0.0014$
RS Obj Mean $25.09$ $69.36$ $684.67$ $6852.08$ $72910.55$
Obj Median $25.35$ $68.58$ $663.10$ $6509.01$ $68904.42$
% Infeasible $0\%$ $3\%$ $0\%$ $1\%$ $39\%$
Time (Sec) $0.0010$ $0.0010$ $0.0012$ $0.0019$ $0.0103$
@@ Appendix G Details for Constraints Violations
In this section, we analyze constraint violations for three benchmark problems. The analysis focuses on both the frequency and magnitude of constraint violations, visualized through heatmaps for a comprehensive understanding. Each heatmap (Figure 6, Figure 7 and Figure 8) illustrates rows as test instances and columns as individual constraints.
@@
The heatmap for the convex quadratic problem (Figure 6) and the simple non-convex problem (Figure 7) reveals a sparse distribution of violations, predominantly concentrated in a single constraint. This indicates that most constraints are consistently satisfied, with only a few isolated violations. Notably, 4 out of 100 test instances of convex quadratic problem under RC are infeasible, and among these, 3 violations occur within the same constraint. Overall, the constraint violations are nearly negligible, confirming the effectiveness of the proposed methods.
@@ Appendix I Training Time Comparison
Table 9, 10 and 11, summarize the training times (in seconds) required by each method for problems of different scales. These results offer a clear view of the computational demands as problem complexity grows.

### v5

@@ Appendix D Neural Network Structure and Hyperparameters
The constraint penalty weight $\lambda$ was set to $300$ for benchmark problems. All networks were trained using the AdamW optimizer with a learning rate of $10^{-3}$ and a batch size of $64$ over $200$ epochs. Early stopping was applied based on validation performance to ensure convergence without overfitting.
@@ Appendix E MINLP Problem Setup and Parameter Sampling
The convex quadratic problems used in our experiments are formulated as follows:
where the coefficients $\bm{Q}\in\mathbb{R}^{n\times n}$, $\bm{p}\in\mathbb{R}^{n}$, and $\bm{A}\in\mathbb{R}^{m\times n}$ were fixed, while $\bm{b}\in\mathbb{R}^{m}$ were treated as parametric coefficients (input features), varying across instances.
where $\bm{Q}\in\mathbb{R}^{n\times n}$ is a diagonal matrix with entries sampled uniformly from $[0,0.01]$, ensuring convexity. The vector $\bm{p}\in\mathbb{R}^{n}$ has entries drawn from a uniform distribution over $[0,0.1]$, while the constraint matrix $\bm{A}\in\mathbb{R}^{m\times n}$ is generated from a normal distribution with a standard deviation of $0.1$. The parameter $\bm{b}\in\mathbb{R}^{m}$, representing the right-hand side of the inequality constraints, is sampled uniformly from $[-1,1]$. These variations in $\bm{b}$ across instances ensure the parametric nature of the problem.
@@
The simple non-convex problem used in the experiments is derived by modifying the convex quadratic programming problem as follows:
where the sine function is applied element-wise to the decision variables $\bm{x}$. This introduces non-convexity into the problem, making it more challenging compared to the convex case. For the simple non-convex problems, the coefficients $\bm{Q}$, $\bm{p}$, $\bm{A}$, and $\bm{b}$ are generated in the same way as in the quadratic formulation. However, an additional parameter $\bm{d}\in\mathbb{R}^{m}$ is introduced, with each element independently sampled from a uniform distribution over $[-0.5,0.5]$. The parameter $\bm{d}$ modifies the constraint matrix $\bm{A}$ by adding $\bm{d}$ to its first column and subtracting $\bm{d}$ from its second column. Alongside $\bm{d}$, the right-hand side vector $\bm{b}$ remains a dynamic parameter in the problem.
@@
The mixed-integer Rosenbrock problem used in this study is defined as:
where $\bm{x}\in\mathbb{R}^{n}$ are continuous decision variables and $\bm{y}\in\mathbb{Z}^{n}$ are integer decision variables. The vectors $\bm{p}\in\mathbb{R}^{n}$ and $\bm{q}\in\mathbb{R}^{n}$ are fixed for each instance, while the parameters $b$ and $\bm{a}$ vary. In details, the vectors $\bm{p}\in\mathbb{R}^{n}$ and $\bm{q}\in\mathbb{R}^{n}$ aregenerated from a standard normal distribution. The parameter $b$ is uniformly distributed over $[1,8]$ for each instance, and the parameter $\mathbf{a}\in\mathbb{R}^{n}$ represents a vector where elements drawn independently from a uniform distribution over $[0.5,4.5]$. The parameters $b$ and $\mathbf{a}$ influence the shape of the feasible region and the landscape of the objective function, serving as input features to the neural network.
@@ Appendix F Ablation Study
The results of the ablation experiments, summarized in Table 5, Table 6 and Table 7, demonstrate the importance of the correction layers $\varphi_{\Theta_{2}}$ in improving both solution quality and feasibility. The experimental setup and model parameters used are consistent with those in the main text, ensuring the results are directly comparable. RL shows a significant drop in feasibility rates, highlighting the importance of incorporating learnable corrective adjustments during training. Similarly, while RS benefits from differentiability via STE, the lack of learnable correction limits its performance compared to RC and LT.
@@
RL Obj Mean $0.563$ $-2.182$ $-4.569$ $-13.732$ $-15.985$ $-37.363$ $-86.385$ $-165.047$
Obj Median $0.390$ $-2.547$ $-4.585$ $-13.755$ $-15.970$ $-37.363$ $-86.385$ $-165.047$
% Infeasible $37\%$ $55\%$ $32\%$ $32\%$ $51\%$ $96\%$ $100\%$ $100\%$
Time (Sec) $0.0005$ $0.0006$ $0.0004$ $0.0004$ $0.0006$ $0.0005$ $0.0006$ $0.0011$
RS Obj Mean $1.087$ $-0.694$ $-3.202$ $-11.213$ $-9.240$ $-21.975$ $-47.732$ $-92.846$
Obj Median $0.926$ $-1.084$ $-3.195$ $-11.244$ $-9.237$ $-21.975$ $-47.732$ $-92.846$
@@
RL Obj Mean $0.374$ $0.741$ $0.054$ $-0.421$ $-0.915$ $-2.177$ $-10.684$ $-21.837$
Obj Median $0.230$ $0.575$ $0.032$ $-0.458$ $-0.915$ $-2.177$ $-10.684$ $-21.837$
% Infeasible $19\%$ $19\%$ $17\%$ $42\%$ $66\%$ $88\%$ $100\%$ $100\%$
Time (Sec) $0.0005$ $0.0005$ $0.0005$ $0.0005$ $0.0007$ $0.0005$ $0.0006$ $0.0012$
RS Obj Mean $0.866$ $1.421$ $0.588$ $2.704$ $3.531$ $7.038$ $15.065$ $37.559$
Obj Median $0.537$ $0.959$ $0.572$ $2.693$ $3.533$ $7.038$ $15.065$ $37.559$
@@
RL Obj Mean / Median $22.19/22.59$ $62.75/62.32$ $609.94/626.64$ $6131.90/5806.59$ $47.00e4/5.55e$ 4
% Infeasible $44\%$ $28\%$ $36\%$ $17\%$ $31\%$
Time (Sec) $0.0005$ $0.0005$ $0.0005$ $0.0006$ $0.0015$
RS Obj Mean / Median $25.26/25.98$ $67.36/65.05$ $667.72/641.89$ $7538.04/7597.98$ $48.41e4/8.44e$ 4
% Infeasible $0\%$ $3\%$ $0\%$ $1\%$ $30\%$
Time (Sec) $0.0010$ $0.0011$ $0.0011$ $0.0021$ $0.0091$
@@ Appendix G Details for Constraints Violations
In this section, we analyze constraint violations for three benchmark problems. The analysis focuses on both the frequency and magnitude of constraint violations, visualized through heatmaps for a comprehensive understanding. Each heatmap (Figure 6, Figure 7 and Figure 8) illustrates rows as test instances and columns as individual constraints. The heatmaps illustrate results for models trained on datasets of 8,000 samples, with penalty weights set to $200$, $100$, and $100$, respectively, for the three problems.
@@
The heatmap for the convex quadratic problem (Figure 6) and the simple non-convex problem (Figure 7) reveals a sparse distribution of violations, predominantly concentrated in a single constraint. This indicates that most constraints are consistently satisfied, with only a few isolated violations. Notably, 4 out of 100 test instances of convex quadratic problem under RC are infeasible, and among these, 3 violations occur within the same constraint, and all 3 infeasible solutions of convex quadratic from LT method also appear in the one constraint. Overall, the constraint violations are nearly negligible, confirming the effectiveness of the proposed methods.
@@ Appendix I Training Time Comparison
Table 9, 10 and 11, summarize the training times (in seconds) required by each method for problems of different scales. These results highlight the computational efficiency of our methods during training, with training times for most problem instances remaining within a few hundred seconds. Even for large-scale problems, such as the 20000×4 Rosenbrock problem, the RC and LT method required only a few hours of training to handle a problem that exact solvers struggle to even find feasible solutions for in reasonable time. This efficiency is largely attributed to our simple neural network architecture, which enables scalable and efficient training. Thus, our method is particularly advantageous in real-world scenarios where rapid deployment and scalability are critical.